Find two consecutive odd numbers whose product exceeds its sum by 119.
Answer by checkley75(3666)

x(x+2)=(x+x+2)+119
x^2+2x=2x+121
x^2-121=0
(x-11)(x+11)
x-11=0
x=11 answer for the smaller odd number.
11+2=13 for the larger odd number.
proof
11*13=11+13+119
143=143